# PriPHiT: Privacy-Preserving Hierarchical Training of Deep Neural Networks

### v1

@@ IV-C Inference Phase and Attacking during the Inference
After finishing the training of the user’s cloud analyzer which is concurrently done with the untrusted cloud’s training of the attackers, the user starts the inference phase. We assume that the inference phase is done on a new dataset that is unseen by all models (Fig. 1-bottom illustrates the scenario). We assess the efficacy of the cloud analyzer and the edge feature extractor in keeping the desired content by studying the accuracy of the desired content classification. We also test the effectiveness of the edge feature extractor to conceal the sensitive content by studying the accuracy of the trained classification attacker on sensitive content and the similarity metrics between the reconstructed images made by the trained reconstruction attacker and the original inputs. Additionally, we consider another scenario in this phase where the parameters of the trained edge feature extractor are leaked. We assume that the untrusted cloud performs a white-box reconstruction attack using these parameters. For this attack, we use a similar method to the one used in the recent work [13]. The details of the white-box reconstruction attack are described in Supplementary Material VIII. We conduct these attacks on all feature maps extracted from the test set and compare them with the original unseen input images quantitatively and qualitatively.
TABLE V: PSNR and SSIM metrics averaged on the output of the deep reconstruction and white-box reconstruction attacks in comparison to the unseen inputs from the test set for different architectures and pairs of attributes in the experiments on the CelebA dataset.
@@
TABLE VI: PSNR and SSIM metrics averaged on the output of the deep reconstruction and white-box reconstruction attacks in comparison to the unseen inputs from the test subset of the FFHQ dataset for the experiment of smiling versus gender using different architectures.
@@
Table III shows the accuracy of the VGG-11, ResNet-18, and MobileViT-xxs architectures trained with PriPHiT and the baselines on the unseen test set. The cloud analyzer accuracy in classifying the desired content and the classification attacker accuracy on the sensitive content are shown for two experiments with different pairs of attributes on the CelebA dataset. The analyzer accuracy in the PriPhiT method is close to the baselines, which represents the pinnacle of achievable accuracy. On the other hand, while the baselines exhibit notable privacy leakage, evidenced by the high accuracy of the classification attack on the sensitive content, PriPHiT provides strong preservation of privacy which results in low accuracy in such attacks. When the privacy budget is lower ($\epsilon=0.5$), PriPHiT becomes stronger in removing the sensitive content which costs a slight decrease in terms of the analyzer accuracy in comparison to higher budgets ($\epsilon=2$) [^1] The effect of $\epsilon$ on the performance of the method is illustrated in Supplementary Material IX. Since the CelebA dataset is not balanced on all attributes, to show the lowest possible accuracy for each experiment, a trivial classifier is also implemented that blindly outputs the majority label. Using the trivial classifier as the minimum and the baseline as the maximum, the effectiveness of the model can be evaluated. For example, in the smiling versus gender experiment using the PriPHiT ($\epsilon=0.5$) method on ResNet-18, PriPHiT can reduce the attacker accuracy down to $61.71\%$ which is just $0.4\%$ higher than the blind trivial classifier. This happens while the analyzer accuracy is $88.94\%$, just $2.14\%$ lower than the maximum accuracy provided by the baseline. Similar accuracy results for the experiment of smiling versus gender on the unseen test subset of the FFHQ dataset are shown in Table IV. Again, it is shown that the PriPHiT method is able to reduce the classification attacker’s accuracy on the sensitive content close to the performance level of the blind trivial classifier. Notably, this achievement is attained with only marginal degradation in the analyzer accuracy. Similar to the experiments on the CelebA dataset, when the privacy budget is low (e.g., $\epsilon=0.5$), the attacker’s accuracy is reduced more, indicating stronger preservation of privacy, while it costs in lower accuracy of the analyzer.
@@
Table V shows the PSNR and SSIM metrics on the outputs of the two reconstruction attacks on the PriPHiT method in comparison to the baselines, averaged on the whole test subset for the experiments on the CelebA dataset. Both metrics reduce significantly using the PriPHiT training method while they show high values on the baselines, which indicate successful defense against reconstruction attacks in PriPHiT. Again, lower privacy budgets ($\epsilon=0.5$) provide better defense in comparison to the higher ones ($\epsilon=2$). Examples of the outputs of the reconstruction attacks together with the unseen inputs for the smiling versus gender experiment are shown in Figs. 5 and 6. We see that the output of both attacks can reconstruct the inputs with gender indicators clearly from the extracted feature maps by the baselines; however, they fail to reconstruct the inputs when the PriPHiT ($\epsilon=1$) method is used. It is interesting to see that in the outputs of the deep reconstruction attacks, the gender indicators (e.g., beard, long hair) are removed while white-box attacks fail completely. Similar figures for the other experiments are shown in Supplementary Material X. Moreover, similar results for the experiments on the FFHQ dataset are observed, as depicted in Table VI. Again our results show a notable decrease in the averaged similarity metrics when employing the PriPHiT method, as compared to the baselines. Smaller privacy budgets (e.g., $\epsilon=0.5$) again result in a stronger preservation of privacy. Examples of the outputs of reconstruction attacks for this experiment are shown in Fig. 7. It is shown that the deep reconstruction and white-box reconstruction attacks can reconstruct the inputs with gender indicators from the extracted feature maps by the baselines; however, they fail when the PriPHiT ($\epsilon=1$) method is employed.
@@ V Conclusion and Future Works
VI Figures of the Three Architectures Used in the Experiments
@@
Figs. 8, 9, and 10 show the details of the VGG-11, ResNet-18, and MobileViT-xxs architectures that have been used in our experiments in the different steps of PriPHiT execution.
@@
VII Results of the Simultaneous Attacks in Mouth Open versus Wearing Makeup Experiment
@@
Figs. 11 and  12 show the results of the simultaneous attacks during the training phase in the experiments when having an open mouth is selected as the desired content and wearing makeup is the sensitive content. Again, there is a substantial privacy leakage in the baselines as the simultaneous classification attack can provide a high accuracy and the deep reconstruction attack can provide a low MSE loss for the reconstructions in comparison to the original inputs of the user. However, in the PriPHiT method, the classification attack accuracies stay low and the MSE losses of the reconstruction attack stay high, indicating unsuccessful attacks. Additionally, the plots reach a plateau and cannot improve with more epochs, since the adversarial privacy-preserving training using the edge adversary is continuously done in this stage. A lower privacy budget ($\epsilon=0.5$) again provides a stronger preservation of privacy in comparison to a higher budget ($\epsilon=2$).
VIII White-Box Reconstruction Attack

### v2

@@ IV-C Inference Phase and Attacking during the Inference
After finishing the training of the user’s cloud analyzer which is concurrently done with the untrusted cloud’s training of the attackers, the user starts the inference phase. We assume that the inference phase is done on a new dataset that is unseen by all models (Fig. 1-bottom illustrates the scenario). We assess the efficacy of the cloud analyzer and the edge feature extractor in keeping the desired content by studying the accuracy of the desired content classification. We also test the effectiveness of the edge feature extractor to conceal the sensitive content by studying the accuracy of the trained classification attacker on sensitive content and the similarity metrics between the reconstructed images made by the trained reconstruction attacker and the original inputs. Additionally, we consider another scenario in this phase where the parameters of the trained edge feature extractor are leaked. We assume that the untrusted cloud performs a white-box reconstruction attack using these parameters. For this attack, we use a similar method to the one used in the recent work [13]. The details of the white-box reconstruction attack are described in Supplementary Material D. We conduct these attacks on all feature maps extracted from the test set and compare them with the original unseen input images quantitatively and qualitatively.
TABLE VII: PSNR and SSIM metrics averaged on the output of the deep reconstruction and white-box reconstruction attacks in comparison to the unseen inputs from the test set for different architectures and pairs of attributes in the experiments on the CelebA dataset.
@@
TABLE VIII: PSNR and SSIM metrics averaged on the output of the deep reconstruction and white-box reconstruction attacks in comparison to the unseen inputs from the test subset of the FFHQ dataset for the experiment of smiling versus gender using different architectures.
@@
Table V shows the accuracy of the VGG-11, ResNet-18, and MobileViT-xxs architectures trained with PriPHiT and the baselines on the unseen test set. The cloud analyzer accuracy in classifying the desired content and the classification attacker accuracy on the sensitive content are shown for two experiments with different pairs of attributes on the CelebA dataset. The analyzer accuracy in the PriPhiT method is close to the baselines, which represents the pinnacle of achievable accuracy. On the other hand, while the baselines exhibit notable privacy leakage, evidenced by the high accuracy of the classification attack on the sensitive content, PriPHiT provides strong preservation of privacy which results in low accuracy in such attacks. When the privacy budget is lower ($\epsilon=0.5$), PriPHiT becomes stronger in removing the sensitive content which costs a slight decrease in terms of the analyzer accuracy in comparison to higher budgets ($\epsilon=2$) [^1] The effect of $\epsilon$ on the performance of the method is illustrated in Supplementary Material E. Since the CelebA dataset is not balanced on all attributes, to show the lowest possible accuracy for each experiment, a trivial classifier is also implemented that blindly outputs the majority label. Using the trivial classifier as the minimum and the baseline as the maximum, the effectiveness of the model can be evaluated. For example, in the smiling versus gender experiment using the PriPHiT ($\epsilon=0.5$) method on ResNet-18, PriPHiT can reduce the attacker accuracy down to $61.71\%$ which is just $0.4\%$ higher than the blind trivial classifier. This happens while the analyzer accuracy is $88.94\%$, just $2.14\%$ lower than the maximum accuracy provided by the baseline. Similar accuracy results for the experiment of smiling versus gender on the unseen test subset of the FFHQ dataset are shown in Table VI. Again, it is shown that the PriPHiT method is able to reduce the classification attacker’s accuracy on the sensitive content close to the performance level of the blind trivial classifier. Notably, this achievement is attained with only marginal degradation in the analyzer accuracy. Similar to the experiments on the CelebA dataset, when the privacy budget is low (e.g., $\epsilon=0.5$), the attacker’s accuracy is reduced more, indicating stronger preservation of privacy, while it costs in lower accuracy of the analyzer.
@@
Table VII shows the PSNR and SSIM metrics on the outputs of the two reconstruction attacks on the PriPHiT method in comparison to the baselines, averaged on the whole test subset for the experiments on the CelebA dataset. Both metrics reduce significantly using the PriPHiT training method while they show high values on the baselines, which indicate successful defense against reconstruction attacks in PriPHiT. Again, lower privacy budgets ($\epsilon=0.5$) provide better defense in comparison to the higher ones ($\epsilon=2$). Examples of the outputs of the reconstruction attacks together with the unseen inputs for the smiling versus gender experiment are shown in Figs. 5 and 6. We see that the output of both attacks can reconstruct the inputs with gender indicators clearly from the extracted feature maps by the baselines; however, they fail to reconstruct the inputs when the PriPHiT ($\epsilon=1$) method is used. It is interesting to see that in the outputs of the deep reconstruction attacks, the gender indicators (e.g., beard, long hair) are removed while white-box attacks fail completely. Similar figures for the other experiments are shown in Supplementary Material F. Moreover, similar results for the experiments on the FFHQ dataset are observed, as depicted in Table VIII. Again our results show a notable decrease in the averaged similarity metrics when employing the PriPHiT method, as compared to the baselines. Smaller privacy budgets (e.g., $\epsilon=0.5$) again result in a stronger preservation of privacy. Examples of the outputs of reconstruction attacks for this experiment are shown in Fig. 7. It is shown that the deep reconstruction and white-box reconstruction attacks can reconstruct the inputs with gender indicators from the extracted feature maps by the baselines; however, they fail when the PriPHiT ($\epsilon=1$) method is employed.
@@
To demonstrate the effectiveness of our privacy-preserving method against more advanced deep reconstruction attacks, we evaluate a more sophisticated attack by training a DeblurGAN-v2 architecture [35] on top of our reconstruction attacker, which was designed based on the suggestions of [8]. The results of this experiment, conducted on the CelebA dataset, are presented in Table IX and Figure 8. As indicated by the results, this reconstruction attack leads to sharper and more precise outputs for the baseline models across different architectures, as evidenced by higher PSNR values in Table IX and clearer visual results in Figure 8. However, when the PriPHiT method is applied, the attack fails to improve the quality of the reconstructed images, highlighting the robustness of our method against GAN-enhanced attacks.
IV-D Experiments on Medical Data
To further demonstrate the effectiveness of our privacy-preserving hierarchical training method on private data types beyond facial images, we conducted experiments on medical data samples. Specifically, we trained similar deep neural network architectures using PriPHiT on the CheXpert dataset [36], which consists of chest X-ray images of patients. For these experiments, we used a subset of CheXpert provided by [37]. We selected having symptoms as the desired attribute and the patient’s gender as the sensitive attribute. Due to a significant class imbalance in the dataset—where symptomatic samples far outnumber benign samples—we balanced the dataset by randomly subsampling symptomatic images to match the number of benign images. The data was then divided into $20$K samples for training (also used for simultaneous attacks during the training phase), $18$K samples for attacker training, and approximately $7$K samples for testing. Similar to the experiment on the FFHQ dataset, VGG-11 and ResNet-18 architectures were used for the experiments on the CheXpert dataset, as the number of samples was not sufficient for training the MobileViT-xxs architecture.
Figure 9 illustrates the results of simultaneous classification and reconstruction attacks over 50 training epochs when our PriPHiT method is applied, compared to baselines with similar architectures but without privacy-preserving features. In the baseline models, the accuracy of simultaneous classification attacks is high and continues to increase with more epochs, indicating a successful attack on the users’ sensitive content. In contrast, when PriPHiT is used, the accuracy of classification attacks remains low and close to that of a trivial classifier that blindly predicts the majority label, demonstrating PriPHiT’s effectiveness in preserving privacy during the training phase. Furthermore, the MSE Loss between the outputs of simultaneous reconstruction attackers and the inputs in the baseline models approaches zero, indicating successful reconstruction attacks. However, when PriPHiT is employed, the MSE Loss remains high and plateaus, further illustrating PriPHiT’s ability to defend against simultaneous reconstruction attacks.
TABLE IX: Performance of the PriPHIT and baseline methods against the DeblurGAN-V2-boosted deep reconstruction attacker [35] in the smiling versus gender experiment on the CelebA dataset.
Architecture Model Deep Reconstruction Attack DeblurGANv2-Boosted
Deep Reconstruction Attack
PSNR SSIM PSNR SSIM
VGG-11 PriPHIT ($\epsilon$=0.5) 12.64 0.31 10.17 0.38
PriPHIT ($\epsilon$=1) 12.47 0.31 10.75 0.35
Baseline 24.74 0.88 28.29 0.88
ResNet-18 PriPHIT ($\epsilon$=0.5) 11.78 0.29 9.88 0.37
PriPHIT ($\epsilon$=1) 12.12 0.30 10.05 0.36
Baseline 24.89 0.89 27.91 0.90
MobileViT-xxs PriPHIT ($\epsilon$=0.5) 11.74 0.28 10.54 0.34
PriPHIT ($\epsilon$=1) 11.77 0.28 10.33 0.36
Baseline 20.05 0.74 22.62 0.75
Figure 8: Examples of the results of the original Deep Reconstruction Attack (shown as DRA) and DeblurGAN-v2-boosted Deep Reconstruction Attack (shown as DGDRA) on PriPHiT ($\epsilon=1$) in comparison to the similar attacks on the baselines using the unseen inputs from the test subset of the CelebA dataset. Smiling and Gender attributes are selected as the desired and sensitive contents, respectively. Note that the DeblurGAN-v2-Boosted attack makes sharper and more clear reconstruction on Baseline model in comparison to the original Deep Reconstruction Attack, while it fails to improve the results when our PriPHiT method is used.
After completing the training phase, the models are evaluated on the unseen test set. The accuracy results of the analyzer and attacker models during the inference phase are presented in Table X. Similar to the previous experiments, the classification attack accuracy on PriPHiT is significantly lower compared to the baseline models, indicating the effectiveness of our method in removing sensitive content from the shared feature map. This accuracy is very close to that of a trivial classifier. Conversely, the accuracy of the analyzer model on the desired content remains high and comparable to the baseline when PriPHiT is applied, demonstrating that PriPHiT preserves task-relevant information. Notably, in some experiments, the analyzer accuracy slightly exceeds that of the baseline, which may be attributed to a regularization effect introduced by the added noise [38].
The similarity metrics for the outputs of reconstruction attacks are shown in Table XI, with examples of the outputs from deep reconstruction and white-box reconstruction attacks provided in Figure 10. As shown in the table, the similarity metrics are significantly lower when the PriPHiT method is applied, particularly with tighter privacy budget values ($\epsilon=0.5$), compared to the baselines. This demonstrates the effectiveness of our method in defending against these attacks. In the baseline models, deep reconstruction attackers successfully generate clear reconstructions of the original unseen inputs, including sensitive gender indicators (e.g., breasts). However, when PriPHiT is used, the deep reconstruction attacker fails to reconstruct these sensitive features. Similarly, white-box reconstruction attacks succeed in the baseline models, producing inputs containing sensitive content, but fail entirely when PriPHiT is employed.
Figure 9: The simultaneous attack during the training phase for the having symptom versus gender experiment on the CheXpert dataset. Left: The simultaneous classification attack accuracy on the feature maps coming from the user’s private data at each epoch. Right: MSE loss of the simultaneous reconstruction attack on the feature maps coming from the user’s private data compared with the user’s original inputs at each epoch.
TABLE X: Comparison of the test accuracy of PriPHiT with the baselines for different privacy budgets using different architectures for the experiments on the CheXpert dataset. The results are provided when the task-relevant content is having symptoms and the sensitive content is selected to be gender. For the experiments, a trivial classifier shows the lowest accuracy level. It is similar to a blind classifier that just outputs the class with the majority of labels.
Experiment Architecture Method Analyzer Accuracy Attacker Accuracy
(desired content) (sensitive content)
VGG-11 PriPHiT ($\epsilon=0.5$) $73.38\%$ $59.16\%$
PriPHiT ($\epsilon=1$) $74.58\%$ $59.86\%$
Smiling Baseline $72.22\%$ $92.05\%$
vs. Gender ResNet-18 PriPHiT ($\epsilon=0.5$) $71.20\%$ $59.18\%$
PriPHiT ($\epsilon=1$) $70.33\%$ $59.18\%$
Baseline $70.76\%$ $89.89\%$
Trivial Classifier $50.00\%$ $59.14\%$
TABLE XI: PSNR and SSIM metrics averaged on the output of the deep reconstruction and white-box reconstruction attacks in comparison to the unseen inputs from the test subset of the CheXpert dataset for the experiment of having symptom versus gender using different architectures.
PSNR SSIM PSNR SSIM
VGG-11 PriPHiT ($\epsilon=0.5$) $14.21$ $0.39$ $8.97$ $0.13$
PriPHiT ($\epsilon=1$) $14.81$ $0.43$ $9.16$ $0.15$
Smiling Baseline $25.59$ $0.92$ $14.02$ $0.55$
vs. Gender ResNet-18 PriPHiT ($\epsilon=0.5$) $13.99$ $0.37$ $9.23$ $0.13$
PriPHiT ($\epsilon=1$) $14.45$ $0.40$ $9.17$ $0.16$
Baseline $21.59$ $0.92$ $13.54$ $0.46$
Figure 10: Examples of the results of the deep reconstruction and white-box reconstruction attacks on PriPHiT ($\epsilon=1$) in comparison to the similar attacks on the baselines using the unseen inputs from the test subset of the CheXpert dataset. Having Symptom and Gender attributes are selected as the desired and sensitive contents, respectively. Note the suppression of gender indicators (e.g., breasts) in the results of deep reconstruction attacks and the failure of white-box reconstruction attacks when the PriPHiT method is used.
@@ V Conclusion and Future Works
A Complexity Analysis of PriPHiT
In this section, we perform a simplified complexity analysis of the PriPHiT method. For simplicity, we assume that the communication between edge and cloud takes a negligible time. We divide this analysis into the two phases of edge pre-training and edge-cloud training (see Sections III-A and III-B).
1. Edge Pre-Training (Section III-A)
  • Feature Extractor: Feature extraction involves forward passes and backward passes through the edge feature extractor, which typically has complexity:
    $O(m\cdot n^{2}\cdot k^{2}),$ (4)
    where $m$ is the number of filters, $n$ is the feature map size, and $k$ is the kernel size. Adding Laplacian noise is $O(n^{2})$ which is not dominant.
  • Analyzer: It contains A forward pass through the analyzer layer (linear/convolutional layer) with complexity $O(d)$, where $d$ is the feature map size with a backward pass for gradient updates which has compelxity $O(d)$. Thus, the total complexity is $O(d)$ per epoch.
  • Adversary: For $m_{a}$ adversary training steps per epoch, each step involves a forward pass through the adversary (linear/convolutional layer) with complexity $O(d)$, where $d$ is the feature map size and a backward pass for gradient updates that has complexity $O(d)$. Thus, the total complexity is $m_{a}\cdot O(d)$ per epoch.
  • Overall Complexity for Edge Pre-Training: With $n_{p}$ epochs and an edge device with a computational speed of $s_{e}$, the total complexity is:
    $O(\frac{n_{p}}{s_{e}}\cdot(m\cdot n^{2}\cdot k^{2}+(m_{a}+1)\cdot d)).$ (5)
2. Edge-Cloud Training (Section III-B)
  • Feature Extraction: Similar to the edge feature extraction, forward and backward passes scale as:
    $O(m\cdot n^{2}\cdot k^{2}+m^{\prime}\cdot n^{\prime 2}\cdot k^{\prime 2}),$ (6)
    where $m^{\prime},n^{\prime},k^{\prime}$ are cloud feature extractor parameters (which exist inside the large cloud analyzer shown in Figure 1).
  • Analyzer: It involves a forward pass through the analyzer decision making layer (linear layer) that has complexity $O(d^{\prime})$, where $d^{\prime}$ is the feature map size and a backward pass for gradient updates with complexity $O(d^{\prime})$. Thus, the total complexity is $O(d^{\prime})$ per epoch.
  • Adversary: Similar to the edge pre-training, the complexity is $m_{a}\cdot O(d)$ per epoch.
  • Overall Complexity for Edge-Cloud Training: Including $n_{t}$ epochs and computational speed of $s_{e}$ and $s_{c}$ for the edge and cloud devices, we have:
    $O(\frac{n_{t}}{s_{e}}\cdot(m\cdot n^{2}\cdot k^{2}+m_{a}\cdot d)+\frac{n_{t}}{s_{c}}\cdot(m^{\prime}\cdot n^{\prime 2}\cdot k^{\prime 2}+d^{\prime})).$ (7)
3. Total Complexity
Combining both stages, the total complexity is:
$O(\frac{n_{p}}{s_{e}}\cdot(m\cdot n^{2}\cdot k^{2}+(m_{a}+1)\cdot d))\ +$ (8)
$O(\frac{n_{t}}{s_{e}}\cdot(m\cdot n^{2}\cdot k^{2}+m_{a}\cdot d)+\frac{n_{t}}{s_{c}}\cdot(m^{\prime}\cdot n^{\prime 2}\cdot k^{\prime 2}+d^{\prime})).$
4. Special Cases
  • Case 1 (Slow Edge Device): If $s_{e}$ is small or, $n_{p}$ is large (slow edge device or too many edge pre-training epochs), edge pre-training dominates the complexity.
  • Case 2 (Large Cloud Models): If $m^{\prime}$, $n^{\prime}$, $k^{\prime}$, or $n_{t}$ grow (large cloud models or too many edge-cloud training epochs), edge-cloud training dominates the complexity.
5. Summary
The PriPHiT method is designed to be computationally efficient, particularly for resource-constrained edge devices, by minimizing parameter counts on the edge. The overall complexity is influenced by key factors, including the computational speed of the edge device ($s_{e}$), model dimensions (parameters $m,n,k,m^{\prime},n^{\prime},k^{\prime}$), and the number of adversarial or training steps. Achieving runtime efficiency requires a careful balance of the pre-training epochs ($n_{p}$), training epochs ($n_{t}$), and adversarial steps ($m_{a}$). In practical applications, as demonstrated in Section IV, the number of edge pre-training steps ($n_{p}$) is typically kept small to mitigate computational demands on resource-constrained edge devices.
B Figures of the Three Architectures Used in the Experiments
@@
Figs. 11, 12, and 13 show the details of the VGG-11, ResNet-18, and MobileViT-xxs architectures that have been used in our experiments in the different steps of PriPHiT execution.
@@
C Results of the Simultaneous Attacks in Mouth Open versus Wearing Makeup Experiment
@@
Figs. 14 and  15 show the results of the simultaneous attacks during the training phase in the experiments when having an open mouth is selected as the desired content and wearing makeup is the sensitive content. Again, there is a substantial privacy leakage in the baselines as the simultaneous classification attack can provide a high accuracy and the deep reconstruction attack can provide a low MSE loss for the reconstructions in comparison to the original inputs of the user. However, in the PriPHiT method, the classification attack accuracies stay low and the MSE losses of the reconstruction attack stay high, indicating unsuccessful attacks. Additionally, the plots reach a plateau and cannot improve with more epochs, since the adversarial privacy-preserving training using the edge adversary is continuously done in this stage. A lower privacy budget ($\epsilon=0.5$) again provides a stronger preservation of privacy in comparison to a higher budget ($\epsilon=2$).
D White-Box Reconstruction Attack